regina says that (3,0) is a solution to the system y= |x-3| and 3x+3y=27 because it satisfies the first equation. rashel says (6,3) a solution to the system because it satisfies. both equations. who is correct? justify the answer.
Answer by jorel555(1290)

Try using the values in the equations:
3(3)+3(0)does NOT equal 27, as per the second equation
3(6)+3(3)=18+9 equals 27.
6-3=3
Thus, rashel(?) is correct, because it satisfies both equations!!!!!!